### v2

# arXiv:2602.22038v2[math.AP] 27 Feb 2026

Abstract
We derive quantitative estimates for the stochastic 2D vortex model on the whole Euclidean space from stochastic particle systems driven by individual and environmental noises, obtaining pathwise quantitative bounds in the terms of relative entropy. The main novelties lie in the application of the Donsker–Varadhan inequality within the framework of moderately interacting particle systems to handle the nonlinear terms, as well as in the introduction of localization techniques combined with a probabilistic data setting, in order to derive estimates for the quadratic variation terms. Moreover, we prove the existence of a suitable solution to the aforementioned model.
MSC2010 subject classification: 49N90, 60H30, 60K35
1 Introduction
In this paper, we establish a quantitative derivation of the stochastic 2D vortex model in the whole Euclidean space, which is a particular case of the following stochastic Fokker–Planck equation
$\mathrm{d}\rho_{t}=\Delta\rho_{t}\,dt+\frac{1}{2}D^{2}\rho_{t}(\sigma\sigma^{\top})_{t}\,dt-\nabla\cdot(\rho_{t}(K\ast\rho_{t}))\,dt-\nabla\rho_{t}\cdot\sigma_{t}\,dB_{t},$ (1)
from a stochastic moderately interacting particle system given by
$dX_{t}^{i,N}$ $=\frac{1}{N}\sum_{k=1}^{N}\left(K\ast V^{N}\right)\left(X_{t}^{i,N}-X_{t}^{k,N}\right)\,dt+\sqrt{2}\,dW_{t}^{i,N}+\sigma_{t}\,dB_{t},\quad$ (2)
where $W_{t}^{i,N}$ and $B_{t}$ are independent standard $\mathbb{R}^{d}$-valued Brownian motions, defined on a filtered probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathbb{P})$, $K$ is a singular kernel and $V^{N}$ a suitable scalling.
To the best of the author’s knowledge, this is the first time that (1) is derived for 2D stochastic vortex model, by moderately interacting particles, in the whole space - extending what was known before for bounded kernels in [32], on $\mathbb{R}^{d}$, and in [39], for sub-Coulombian kernels, under periodic boundary conditions - with pathwise quantitative bounds, in the sense of the relative entropy functional.
In the particle system (2), the stochastic term $\sigma_{t}dB_{t}$ accounts for a source of randomness that affects all particles, modeling a shared environmental fluctuation. Models of this nature appear in several contexts, such as problems in statistical mechanics, population dynamics and biology, as well as in the analysis of numerical Monte–Carlo schemes and related applications; see, e.g., [4], [5], [24], [19], and [29].
The main goal of this paper is to investigate the asymptotics in $N$ of the empirical measure defined by
$S_{t}^{N}\doteq\frac{1}{N}\sum_{i=1}^{N}\delta_{X_{t}^{i,N}},\qquad t\geqslant 0,$ (3)
where $\delta_{a}$ is the delta Dirac measure concentrated at $a\in\mathbb{R}^{d}$. For that purpose, considering $V^{N}\doteq N^{\beta}V(N^{\frac{\beta}{d}}\cdot)$, we introduce the mollified empirical measure
$\rho_{t}^{N}\doteq V^{N}\ast S_{t}^{N}=\int_{\mathbb{R}^{d}}V^{N}(\cdot-y)\,S_{t}^{N}(dy),$
in order to apply the relative entropy functional directly to the densities $\rho^{N}$ and $\rho$, since by the Csiszar-Kullback-Pinsker inequality, we get
$\|S^{N}_{t}-\rho_{t}\|_{0}^{2}\lesssim N^{-\frac{\beta}{d}}+\|\rho^{N}_{t}-\rho_{t}\|_{1}^{2}\overset{\text{CKP}}{\lesssim}N^{-\frac{\beta}{d}}+\mathcal{H}\left(\rho^{N}_{t}|\rho_{t}\right),$ (4)
where the metric in the left hand side of (4), is the Kantorovich-Rubinstein metric between two probatility measures. It follows that, if there exists a constant $C=C_{\omega}>0$ such that for any $N\in\mathbb{N}$,
$\sup_{t\in[0,T]}\mathcal{H}(\rho_{t}^{N}|\rho_{t})\leq C(\mathcal{H}(\rho_{0}^{N}|\rho_{0})+N^{-\theta}),$ (5)
where $\theta$ is an explicit positive parameter, by (4) and (5), we have, in particular, propagation of chaos for the marginals of the empirical measure of the particle system, see [1, Section 8.3] and [45].
In the present work, we apply for the first time in the context of moderately interacting particles, the Donsker–Varadhan inequality together with the dissipation provided by the Fisher information, to handle the nonlinearity. In addition, we employ localization techniques combined with the probabilistic data setting, to derive estimates for the quadratic variation terms in the whole space, as the estimate established in [39], under periodic boundary conditions, does not apply in the present framework. Moreover, we prove the existence of a suitable solution to the equation (1).
Related works
Quantitative estimates for singular interactions have been substantially developed in recent years. The relative entropy method for establishing quantitative propagation of chaos in McKean–Vlasov systems with singular interaction kernels was first introduced in [23]. It was developed in the context of general first-order systems with interaction kernels in $W^{-1,\infty}$, a framework that includes the point vortex model approximating the two-dimensional Navier–Stokes equation. We also refer to [42], where the mean-field limit and quantitative propagation of chaos for McKean–Vlasov equations with singular interaction kernels are established via the modulated energy method. In contrast to the relative entropy approach, which operates at the level of the joint law of the particle system, the modulated energy method is formulated directly in terms of the empirical measure associated with the particles.
Since then, the aforementioned works have paved the way for establishing quantitative propagation of chaos results. In [2] and [3], the strategies given in [23] and [42], were adapted to get quantitative derivation for the Patlak–Keller–Segel model, in optimal subcritical regimes. Additionally, [20], by means of a logarithmic Sobolev inequality, derived uniform in time propagation of chaos for the 2D vortex model, adapting the strategy in [23]. Alternatively, with sharp estimates, global in time derivation of that model, based on relaxation estimates for the limiting equation and inequalities of Riesz modulated energy, was derived by [12], see also [41],[11] and [31], for the modulated energy method. Regarding sub-Coulombian kernels, [6] obtained the derivation of the mean-field approximation for Landau-like equations, while [40] obtained global-in-time mean-field convergence for gradient (conservative) diffusive flows. By extending the BBGKY hierarchy given in [30], for singular kernels, [47] obtained the optimal convergence rate for the relative entropy between the k-particle marginals and the corresponding tensorized law for weakly interacting diffusions with bounded interaction kernels.
With this in view we observe that, the relative entropy method has proven effective for deriving quantitative estimates, for singular kernels, under periodic boundary conditions, while the modulated energy method, which can be applied in the whole space, has primarily been developed for deterministic interaction. However, more recently, quantitative propagation of chaos for the 2D vortex model and the 2D Log Gas on the whole space was obtained in [48] and [49], [7], respectively. In these works, the authors exploit Li–Yau-type estimates and Hamilton-type heat kernel estimates to derive suitable bounds for the limiting solution and its derivatives up to second order, which allows them to extend the strategy given in [23], to $\mathbb{R}^{d}$.
Quantitative estimates in the whole space were derived, in the context of moderately interacting particles  [^2] The moderate interact model was introduced by Oelschlager in [34],[33] and [35]. with singular kernels - including the Biot-Savart, the Keller-Segel and the Dirac delta measure kernels - by [36] and [37], based on a semigroup approach introduced in [17], with further advancements in [8], [21], [38], and [27]. It is important to note that these works do not rely on the relative entropy method, but instead establish quantitative estimates in strong functional topologies. The connection between the time evolution of the relative entropy, of the joint distribution, and moderate interacting regime, on the whole space was carried out by [10], based on a combination between the relative entropy and the regularised $L^{2}$-estimate in [35], deriving a propagation of chaos result for the viscous porous medium equation.
But was in [39], on the torus, that the equation (1), was derived, for singular kernels and common noise, in the context of moderate interaction regime, by applying the relative entropy directly to the mollified empirical measure. Earlier, also the same domain, the results established in [36] were extended to the setting with environmental noise in [26]. It is worth highlighting, that quantitative entropy estimates for systems with individual and common noise, as shown in [43] for incompressible the Navier-Stokes equations, [9] for the Hegselmann-Krause model, and [32] for mean-field systems with bounded kernels. For more results, see [18], [15], [16], and [28].
In contrast to the aforementioned works, our application of the relative entropy method provides pathwise estimates, deriving bounds directly at the trajectory level of the particle system. In those works, the estimates are obtained at the level of the joint law.
In order to present our results, this work is organized as follows. In the next two subsections, 1.1 and 1.2, respectively, we introduce the notations and assumptions used throughout the text, setting the stage for the subsequent discussions. In the subsection 1.3, we present our main results. The section 2 is dedicated to the proof of our main results. To ease the presentation, the Appendix, provides the corresponding suitable notion of a solution for the limiting equation, as well as some inequalities used throughout the proofs.
1.1 Notations
For $d\geq 1$, the space of probability density functions on $\mathbb{R}^{d}$ is denoted by $\mathcal{P}(\mathbb{R}^{d})$. For a normed vector space $U$, we denote its norm by $\|\cdot\|_{U}$, except for $L^{p}$ spaces, where we write $\|\cdot\|_{p}$, for $p\geq 1$. The Kantorovich-Rubinstein metric reads, for any two probability measures $\mu$ and $\nu$ on $\mathbb{R}^{d}$,
$\|\mu-\nu\|_{0}\doteq\sup\left\{\int_{\mathbb{R}^{d}}\phi\,d(\mu-\nu)\,;\|\phi\|_{\infty},\|\phi\|_{\text{Lip}}\leq 1\right\}.$ (6)
Let $f$ and $g$ be positive probability density functions on $\mathbb{R}^{d}$. The relative entropy (or Kullback–Leibler divergence) of $f$ with respect to $g$ is defined as
$\mathcal{H}\left(f|g\right)\doteq\int_{\mathbb{R}^{d}}f(x)\ln{\left(\frac{f(x)}{g(x)}\right)}\,dx.$
Also the Fisher information of $f$ with respect to $g$ is given by
$\mathcal{I}\left(f|g\right)\doteq\int_{\mathbb{R}^{d}}f(x)\left|\nabla\ln{\frac{f(x)}{g(x)}}\right|^{2}\,dx=\int_{\mathbb{R}^{d}}\frac{(g(x))^{2}}{f(x)}\left|\nabla\left(\frac{f(x)}{g(x)}\right)\right|^{2}\,dx.$
1.2 Assumptions
  1. $(\mathbf{A}^{V})$
    Let $V^{N}:\mathbb{R}^{d}\to\mathbb{R}$ be the scaling given by $V^{N}(y)=N^{\beta}V(N^{\frac{\beta}{d}}y)$, $\beta\in(0,1)$, with $V:\mathbb{R}^{d}\to\mathbb{R}$ satisfying
    $\left|\nabla V\right|\leq C_{d}V,$ (7)
    and for $\alpha>d/2$
    $V(y)\leq C_{d}(1+|y|^{2})^{-\alpha}.$ (8)
  2. $(\mathbf{A}^{\nabla\cdot K})$
    It holds $\nabla\cdot K=0$ and there exists $K_{0}\in L^{\infty}$, with $\|K_{0}\|_{\infty}\leq 1/4$, such that
    $K=\nabla\cdot K_{0}.\,$ (9)
  3. $(\mathbf{A}^{\sigma})$
    The coefficient $\sigma:[0,T]\to\mathbb{R}^{d\times d}$ is measurable and bounded.
  4. $(\mathbf{A}^{\rho_{0}})$
    The initial condition $\rho_{0}$ is taken such that $\rho_{0}\in W^{2,\infty}\left(\mathbb{R}^{d}\right)$ and
    $\left|\nabla\ln{\rho_{0}}(x)\right|^{2}\leq\widetilde{C}_{1}(1+|x|^{2}),$ (10)
    $\left|\nabla^{2}\ln{\rho_{0}}(x)\right|\leq\widetilde{C}_{2}(1+|x|^{2})$ (11)
    and
    $\rho_{0}(x)\leq\widetilde{C}_{3}\exp{\left(-\widetilde{C}_{3}^{-1}|x|^{2}\right)}.$ (12)
  5. $(\mathbf{A}^{X_{0}})$
    The initial particle system $(X_{0}^{i,N})_{i=1}^{N}$, $N\in\mathbb{N}$ is a sequence of i.i.d with law $\rho_{0}$.
Remark 1.
For an example of mollifier $V:\mathbb{R}^{d}\to\mathbb{R}$ satisfying $\left(\mathbf{A}^{V}\right)$, take
$V(y)\doteq\overline{C}_{d}\exp{\left(-\sqrt{1+|y|^{2}}\right)}$
with $\overline{C}_{d}\doteq\left(\int_{\mathbb{R}^{d}}\exp{\left(-\sqrt{1+|y|^{2}}\right)}\,dy\right)^{-1}$.
Remark 2.
The Biot-Savart kernel in dimension two is given by
$K(x)=\frac{1}{2\pi}\frac{x^{\perp}}{|x|^{2}}$ (13)
and then, it is divergence in the sense of distributions, of a $L^{\infty}$ matrix $K_{0}$, namely $K=\nabla\cdot K_{0}$, with
$K_{0}(x_{1},x_{2})=\left(\frac{1}{2\pi}\arctan{\left(\frac{x_{1}}{x_{2}}\right)}\right)I.$ (14)
It follows that, this singular kernel satisfies the Assumption $\left(\mathbf{A}^{\nabla\cdot K}\right)$.
Remark 3.
Assumption $(\mathbf{A}^{\rho_{0}})$ was formulated in [7, 48, 49], where, relying on Li–Yau–type gradient bounds and Hamilton–type heat kernel estimates, the authors use it to control the solution and its derivatives up to second order. It allows the relative entropy approach of [23] to be extended from the periodic framework to the whole space.
1.3 Statement of the main results
The first result, whose proof is postponed to the Appendix, establishes that the limiting equation is well posed in the appropriate functional framework. Indeed, it provides the regularity properties required in the proof of the main results.
Theorem 1.
Assume $(\mathbf{A}^{\nabla\cdot K})$, $(\mathbf{A}^{\sigma})$, and $(\mathbf{A}^{\rho_{0}})$. Then, there exist $T_{1}>0$, $C_{3}>0$, such that if $\widetilde{C}_{3}<C_{3}$,
  • there exists a pathwise solution of (1) with initial condition $\rho_{0}$, such that $\rho\in L^{\infty}\left([0,T_{1}];W^{2,1}\cap W^{2,\infty}(\mathbb{R}^{d})\right)$, $\mathbb{P}$-a.s.;
  • the solution $\rho$ verifies, $\mathbb{P}$-a.s., that
    $\left|\nabla\ln{\rho}(t,x)\right|^{2}\leq C_{1}(1+|x-X_{t}|^{2}),$ (15)
    $\left|\nabla^{2}\ln{\rho}(t,x)\right|\leq C_{2}(1+|x-X_{t}|^{2})$ (16)
    and
    $\rho(t,x)\leq 2\exp{(-2\pi|x-X_{t}|^{2})},$ (17)
    where
    $dX_{t}=\sigma_{t}dB_{t},$ (18)
    for all $t\in[0,T_{1}]$ and constants $C_{1},C_{2}$ dependent on $\widetilde{C}_{1}$, $\widetilde{C}_{2}$ and $\widetilde{C}_{3}$.
We are now able to extend what was known before for bounded kernels in the whole space as in [32] and in [39], for sub-Coulombian kernels, under periodic boundary conditions. Our main result can be stated as follows.
Theorem 2.
Assume $(\mathbf{A}^{V})$, $(\mathbf{A}^{\nabla\cdot K})$, $(\mathbf{A}^{\sigma})$, $(\mathbf{A}^{\rho_{0}})$, $(\mathbf{A}^{X_{0}})$ and let $\rho$ be a solution of (1), given by Theorem 1. In addition, let the dynamics of the particle system be given by (2) and
$\mathcal{H}(\rho_{0}^{N}|\rho_{0})\leq C,\,\,\,\mathbb{P}-a.s.$
where
$\theta=\min\left(1-\beta(1+\frac{2}{d}+2\alpha);\frac{1}{2}-\beta\Big(1+\frac{1}{d}\Big)\right)-\delta$
with $\delta>0$, such that $\theta>0$, $d\geq 1$ and $\beta\in\left(0,\frac{1}{2\big[1+\frac{1}{d}\big]}\right)\cap\left(0,\frac{1}{\big[1+\frac{2}{d}+2\alpha\big]}\right)$.
Then, there exist a time $T\in(0,T_{1})$ and a constant $C>0$, such that, almost surely for $N>>1$,
$\sup_{t\in[0,T]}\mathcal{H}(\rho_{t}^{N}|\rho_{t})$ $\lesssim\left(\mathcal{H}(\rho_{0}^{N}|\rho_{0})+N^{-\theta}+A_{0}N^{-\theta}+C\right)$ (19)
As a corollary of the previous theorem, we establish propagation of chaos for the marginals of the empirical measure of the particle system. In the sense of [1, Section 8.3], it corresponds to convergence of the empirical measure with respect to the Kantorovich–Rubinstein metric, given in (6).
Corollary 1.
Under assumptions of Theorem 2, we have
$\sup_{t\in[0,T]}\left\|S_{t}^{N}-\rho_{t}\right\|_{0}^{2}$ $\lesssim\left(\mathcal{H}(\rho_{0}^{N}|\rho_{0})+N^{-\theta}+A_{0}N^{-\theta}+C\right)$ (20)
Remark 4.
We observe that, compared with previous works on moderate particle systems without common noise, we lose the order of convergence. This is because the estimation of the martingale term in our work is of order $\frac{1}{2}-\beta\Big(1+\frac{1}{d}\Big)$, while in [36] is the order $\frac{1}{2}-\frac{\beta}{2}$, see Theorem 1.3 in [36]. When environmental noise is taken into account, the overall framework remains comparable to that of [26]. For the same reason, the range of the parameter $\beta$ is smaller in our work. However, in the present work we obtain an improved convergence rate compared to that established in [39].
2 Proofs of main results
In this section we present the proofs of our main results.
In Subsection 2.1, by applying the Itô’s formula, such as in [39], we derive an evolution equation for the relative entropy functional of the regularized empirical measure with respect to the solution of (1).
In Subsection 2.2, we deal with the quadratic variation terms arising from the application of Itô’s formula.
In Subsection 2.3, we address the nonlinear contributions appearing in the evolution equation for the relative entropy functional obtained in the first step.
Finally, in Subsections 2.4 and 2.5, combining the previous estimates with Gronwall’s lemma and suitable localization techniques, we close the argument and complete the proof.
2.1 Time evolution of the relative entropy
By applying Itô’s formula, following the same arguments developed in Section 2 of [39], we derive the corresponding evolution equation for the entropy functional between the regularized empirical measure $\rho^{N}$ and the solution $\rho$ of (1):
$\mathcal{H}(\rho_{t}^{N}|\rho_{t})-\mathcal{H}(\rho_{0}^{N}|\rho_{0})$ $=I_{t}+II_{t}+III_{t}+M_{t}^{N}+IV_{t},$ (21)
where
$I_{t}$ $\overset{}{=}\int_{0}^{t}\int_{\mathbb{R}^{d}}\rho_{s}^{N}\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)[K\ast\rho_{s}-K\ast\rho_{s}^{N}]\,dx\,ds$
$-\int_{0}^{t}\int_{\mathbb{R}^{d}}\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\left\langle S_{s}^{N},V^{N}(x-\cdot)[K\ast\rho_{s}^{N}(\cdot)-K\ast\rho_{s}^{N}(x)]\right\rangle\,dx\,ds,$ (22)
$II_{t}+IV_{t}$ $=\frac{1}{2N^{2}}\sum_{i=1}^{N}\int_{\mathbb{R}^{d}}\int_{0}^{t}\frac{1}{\rho_{s}^{N}}|\nabla V^{N}(x-X_{s}^{i,N})|^{2}\,ds\,dx,$ (23)
$III_{t}$ $=-\frac{1}{2}\int_{0}^{t}\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)\,ds,$ (24)
and
$M_{t}^{N}$ $=\int_{\mathbb{R}^{d}}\int_{0}^{t}\left[\ln(\rho_{s})-\ln(\rho_{s}^{N})\right]\sigma^{\top}_{s}\nabla\rho_{s}^{N}\,dB_{s}\,dx$
$+\int_{\mathbb{R}^{d}}\int_{0}^{t}\sigma^{\top}_{s}\nabla\rho_{s}^{N}\,dB_{s}\,dx$
$+\frac{1}{N}\sum_{i=1}^{N}\int_{\mathbb{R}^{d}}\int_{0}^{t}\left[\ln(\rho_{s})-\ln(\rho_{s}^{N})\right]\nabla V^{N}(x-X_{s}^{i})\,dW^{i}_{s}\,dx$
$-\frac{1}{N}\sum_{i=1}^{N}\int_{\mathbb{R}^{d}}\int_{0}^{t}\nabla V^{N}(x-X_{s}^{i})\,dW^{i}_{s}\,dx$
$+\int_{\mathbb{R}^{d}}\int_{0}^{t}\frac{\rho_{s}^{N}}{\rho_{s}}\sigma^{\top}_{s}\nabla\rho_{s}\,dB_{s}\,dx.$ (25)
2.2 Estimates for $II_{t}+IV_{t}$: quadratic variation terms
We now derive estimates for the quadratic variation terms appearing in (21). The main novelty consists in introducing a suitable stopping time in order to handle the integral over the whole space arising in this term. It is important to emphasize that the argument used in [39], which relies on periodic boundary conditions, is no longer available in the present setting.
First, taking into account Assumption $(\mathbf{A}^{V})$ in (23), we have
$II_{t}+IV_{t}$ $\leq\frac{N^{2\beta/d}}{2N^{2}}\sum_{i=1}^{N}\int_{\mathbb{R}^{d}}\int_{0}^{t}\frac{1}{\rho_{s}^{N}}|V^{N}(x-X_{s}^{i,N})|^{2}\,ds\,dx.$ (26)
We introduce the following stopping time:
$\tau^{N}=\inf\left\{t\geq 0;\exists i\in\{1,...,N\},|X_{t}^{i,N}|\geq N^{\beta}\right\}\wedge T.$ (27)
By decay of $\rho_{0}$ in (12), along with Borel-Cantelli Lemma, almost surely, there exists $N_{0}=N_{0}(\omega)$ such that $\tau^{N}(\omega)>0$, for all $N\geq N_{0}$.
Indeed, setting
$\Lambda^{N}\doteq\left\{\exists i\in\{1,...,N\},|X_{0}^{i,N}|\geq N^{\beta}\right\}=\bigcup_{i=1}^{N}\left\{|X_{0}^{i,N}|\geq N^{\beta}\right\}$ (28)
by Assumption $(\mathbf{A}^{X_{0}})$, we derive
[…]
$=\sum_{i=1}^{N}\int_{|x|\geq N^{\beta}}\rho_{0}(x)\,dx$
$\leq C_{d}\sum_{i=1}^{N}(N^{\beta})^{d-2}e^{-\widetilde{C}_{3}^{-1}(N^{\beta})^{2}}$
$\leq C_{d}N(N^{\beta})^{d-2}e^{-\widetilde{C}_{3}^{-1}(N^{\beta})^{2}}$ (29)
where in the second inequality we are using (12) to get
$\int_{|x|\geq N^{\beta}}\rho_{0}(x)dx$ $\leq\widetilde{C}_{3}\int_{|x|\geq N^{\beta}}e^{-\widetilde{C}_{3}^{-1}|x|^{2}}dx$
$=C_{d}\int_{r=N^{\beta}}^{\infty}e^{-\widetilde{C}_{3}^{-1}r^{2}}r^{d-1}dr\leq C_{d}(N^{\beta})^{d-2}e^{-\widetilde{C}_{3}^{-1}(N^{\beta})^{2}}.$ (30)
It follows by Borel-Cantelli Lemma,
$\mathbb{P}\left(\bigcap_{M=1}^{\infty}\bigcup_{N=M}^{\infty}\Lambda^{N}\right)=0.$ (31)
With this in view for all $\omega\in\Lambda_{0}$, with
$\Lambda_{0}\doteq\bigcup_{M=1}^{\infty}\bigcap_{N=M}^{\infty}(\Lambda^{N})^{c}$ (32)
there exists $N_{0}=N_{0}(\omega)$, such that $\tau^{N}(\omega)>0$, for all $N\geq N_{0}$, since $t\mapsto X_{t}^{i,N}(\omega)$ is continuous, almost surely.
Let $s\in(0,\tau^{N})$. We observe that,
$I^{N}\doteq\int_{\mathbb{R}^{d}}\frac{|V^{N}(x-X_{s}^{i,N})|^{2}}{\rho^{N}_{s}(x)}\,dx$ $=\int_{|x|\leq 2|X_{s}^{i,N}|}\frac{|V^{N}(x-X_{s}^{i,N})|^{2}}{\rho^{N}_{s}(x)}\,dx$
$+\int_{|x|>2|X_{s}^{i,N}|}\frac{|V^{N}(x-X_{s}^{i,N})|^{2}}{\rho^{N}_{s}(x)}\,dx.$ (33)
Now, if $|x|\leq 2|X_{s}^{i,N}|$, then $|x|\leq 2N^{\beta}$ for $s\in(0,\tau^{N})$, and since $V^{N}\leq N^{\beta}$, we derive for $\alpha>d/2$
$V^{N}(x-X_{s}^{i,N})\leq N^{\beta}\left(1+|x|^{2}\right)^{\alpha-\alpha}\leq C_{\alpha}N^{2\alpha\beta+\beta}\left(1+|x|^{2}\right)^{-\alpha}.$ (34)
On the other hand, if $|x|>2|X_{s}^{i,N}|$ we have $|x-X_{s}^{i,N}|>\frac{|x|}{2}$ and by decay of $V$, given in (8), we get
[…]
$\leq C_{\alpha}N^{\beta}\left(1+\frac{|N^{\frac{\beta}{d}}x|^{2}}{4}\right)^{-\alpha}\leq C_{\alpha}N^{\beta}\left(1+\frac{|x|^{2}}{4}\right)^{-\alpha}.$ (35)
Therefore by (33), (34) and (35), we find
$I^{N}$ $\leq C_{\alpha}\int_{|x|\leq 2|X_{s}^{i,N}|}\frac{|V^{N}(x-X_{s}^{i,N})|}{\rho^{N}_{s}(x)}\left(N^{2\alpha\beta+\beta}\left(1+|x|^{2}\right)^{-\alpha}\right)\,dx$
$+C_{\alpha}\int_{|x|>2|X_{s}^{i,N}|}\frac{|V^{N}(x-X_{s}^{i,N})|}{\rho^{N}_{s}(x)}\left(N^{\beta}\left(1+\frac{|x|^{2}}{4}\right)^{-\alpha}\right)\,dx$
$\leq C_{\alpha}N^{2\alpha\beta+\beta}\int_{\mathbb{R}^{d}}\frac{|V^{N}(x-X_{s}^{i,N})|}{\rho^{N}_{s}(x)}\left(1+|x|^{2}\right)^{-\alpha}\,dx$
$+C_{\alpha}N^{2\alpha\beta+\beta}\int_{\mathbb{R}^{d}}\frac{|V^{N}(x-X_{s}^{i,N})|}{\rho^{N}_{s}(x)}\left(1+\frac{|x|^{2}}{4}\right)^{-\alpha}\,dx$
$=C_{\alpha}N^{2\alpha\beta+\beta}\int_{\mathbb{R}^{d}}\frac{|V^{N}(x-X_{s}^{i,N})|}{\rho^{N}_{s}(x)}\phi(x)\,dx,$ (36)
with
$\phi(x)\doteq\left(\left(1+|x|^{2}\right)^{-\alpha}+\left(1+\frac{|x|^{2}}{4}\right)^{-\alpha}\right)$ (37)
Hence by (36) in (26) we obtain
[…]
$\overset{}{\leq}C_{\alpha}\frac{N^{\beta(1+\frac{2}{d}+2\alpha)}}{N^{2}}\int_{0}^{t\wedge\tau^{N}}\sum_{i=1}^{N}\int_{\mathbb{R}^{d}}\frac{|V^{N}(x-X_{s}^{i,N})|}{\rho^{N}_{s}(x)}\phi(x)\,dx\,ds$
$\overset{}{\leq}C_{\alpha}\frac{N^{\beta(1+\frac{2}{d}+2\alpha)}}{N}\int_{0}^{t\wedge\tau^{N}}\int_{\mathbb{R}^{d}}\frac{\frac{1}{N}\sum_{i=1}^{N}V^{N}(x-X_{s}^{i,N})}{\rho^{N}_{s}(x)}\phi(x)\,dx\,ds$
$\overset{}{\leq}C_{\alpha}\frac{N^{\beta(1+\frac{2}{d}+2\alpha)}}{N}\int_{0}^{t\wedge\tau^{N}}\int_{\mathbb{R}^{d}}\phi(x)\,dx\,ds$
$\overset{\alpha>d/2}{\leq}C_{\alpha,T}\frac{N^{\beta(1+\frac{2}{d}+2\alpha)}}{N}.$ (38)
2.3 Estimates for $I_{t}$: the nonlinear terms
We now address the nonlinear terms in (21). The main novelty is to use Donsker-Varadhan inequality (82), in the context of moderate interactions, along with the dissipation that comes from of Fisher information and Assumption $(\mathbf{A}^{\nabla\cdot K})$, to tackle with the singularity of kernel $K$.
First, we set
$I_{t}$ $\overset{}{=}\int_{0}^{t}\int_{\mathbb{R}^{d}}\rho_{s}^{N}\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)[K\ast\rho_{s}-K\ast\rho_{s}^{N}]\,dx\,ds$
$-\int_{0}^{t}\int_{\mathbb{R}^{d}}\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\left\langle S_{s}^{N},V^{N}(x-\cdot)[K\ast\rho_{s}^{N}(\cdot)-K\ast\rho_{s}^{N}(x)]\right\rangle\,dx\,ds$
$\doteq I^{1}_{t}-I^{2}_{t},$ (39)
where
$I^{1}_{t}$ $=-\int_{0}^{t}\int_{\mathbb{R}^{d}}\left[\rho_{s}^{N}\frac{\nabla^{2}\rho_{s}}{\rho_{s}}\right][K_{0}\ast\rho_{s}-K_{0}\ast\rho_{s}^{N}]\,dx\,ds$
$-\int_{0}^{t}\int_{\mathbb{R}^{d}}\left[\nabla\rho_{s}\nabla\left(\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right][K_{0}\ast\rho_{s}-K_{0}\ast\rho_{s}^{N}]\,dx\,ds$
$\doteq-I^{1,1}_{t}-I^{1,2}_{t}.$ (40)
Now by $\epsilon$-Young inequality and convolution inequality, we deduce
$I_{t}^{1,2}$ $=\int_{0}^{t}\int_{\mathbb{R}^{d}}\left[\nabla\rho_{s}\left(\frac{\sqrt{\rho_{s}^{N}}}{\rho_{s}}\right)\left(\frac{\rho_{s}}{\sqrt{\rho_{s}^{N}}}\right)\nabla\left(\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right][K_{0}\ast\rho_{s}-K_{0}\ast\rho_{s}^{N}]\,dx\,ds$
$\leq\epsilon\int_{0}^{t}\int_{\mathbb{R}^{d}}\frac{\rho_{s}^{2}}{\rho_{s}^{N}}\left|\nabla\left(\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right|^{2}\,dx\,ds+C_{\epsilon}\int_{0}^{t}\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{s}|^{2}}{\rho_{s}^{2}}\rho_{s}^{N}\left|K_{0}\ast(\rho_{s}-\rho_{s}^{N})\right|^{2}\,dx\,ds$
$\leq\epsilon\int_{0}^{t}\int_{\mathbb{R}^{d}}\frac{\rho_{s}^{2}}{\rho_{s}^{N}}\left|\nabla\left(\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right|^{2}\,dx\,ds+C_{\epsilon}\int_{0}^{t}\|K_{0}\|\|\rho_{s}-\rho_{s}^{N}\|_{1}^{2}\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{s}|^{2}}{\rho_{s}^{2}}\rho_{s}^{N}\,dx\,ds.$ (41)
We observe that by Donsker-Varadhan inequality (82), for $\eta>>1$
[…]
$+\eta\ln{\left(\int_{\mathbb{R}^{d}}\exp{(-2\pi|x-X_{t}|^{2})}\exp{\left(\frac{C(1+|x-X_{t}|^{2})}{\eta}\right)}\,dx\right)}$
$\overset{}{\leq}\eta\mathcal{H}(\rho_{s}^{N}|\rho_{s})+C_{\eta},$
which implies by Csiszár-Kullback-Pinsker inequality (81) and (41)
$I_{t}^{1,2}$ $\leq\epsilon\int_{0}^{t}\mathcal{I}(\rho^{N}_{s}|\rho_{s})\,ds+C_{\epsilon,\eta,K_{0}}\int_{0}^{t}\mathcal{H}(\rho^{N}_{s}|\rho_{s})\,ds.$ (42)
Regarding term $I_{t}^{1,1}$ in (40), by convolution inequality, we find
$I_{t}^{1,1}$ $=\int_{0}^{t}\|K_{0}\|_{\infty}\|\rho^{N}_{s}-\rho_{s}\|_{1}\int_{\mathbb{R}^{d}}\left|\frac{\nabla^{2}\rho_{s}}{\rho_{s}}\right|\rho^{N}_{s}\,dx\,ds.$ (43)
In addition, if $C_{1,2}\doteq\max{(C_{1},C_{2})}$ since
[…]
by Donsker-Varadhan inequality (82), we obtain
[…]
$+\eta\ln{\left(\int_{\mathbb{R}^{d}}2\exp{(-2\pi|x-X_{t}|^{2})}\exp{\left(\frac{C_{1,2}(1+|x-X_{t}|^{2})}{\eta}\right)}\,dx\right)}$
$\overset{}{\leq}\eta\mathcal{H}(\rho_{s}^{N}|\rho_{s})+C_{\eta},$ (45)
for $\eta>>1$.
Therefore, by (45) in (43), we get
$I_{t}^{1,1}$ $\overset{}{\leq}C_{\eta,K_{0}}\int_{0}^{t}\mathcal{H}(\rho_{s}^{N}|\rho_{s})\,ds+TC_{\eta}.$ (46)
[…]
$I_{t}^{1}$ $\leq C_{\epsilon,K_{0},\eta}\int_{0}^{t}\mathcal{H}(\rho_{s}^{N}|\rho_{s})\,ds+\epsilon\int_{0}^{t}\mathcal{I}(\rho_{s}^{N}|\rho_{s})\,ds+TC_{\eta}.$ (47)
Now, we will focus on the term $I_{t}^{2}$ in (39). First, we recall that by Assumption $\left(\mathbf{A}^{\nabla\cdot K}\right)$, $K=\nabla\cdot K_{0}$ with $K_{0}\in L^{\infty}$. Hence, we arrive at
$\left|K\ast\rho_{s}^{N}(\cdot)-K\ast\rho_{s}^{N}(x)\right|$ $=\left|K_{0}\ast\nabla\rho_{s}^{N}(\cdot)-K_{0}\ast\nabla\rho_{s}^{N}(x)\right|$
$\leq\int_{\mathbb{R}^{d}}\left|K_{0}(\cdot-y)-K_{0}(x-y)\right||\nabla\rho_{s}^{N}(y)|\,dy$
$\leq\int_{\mathbb{R}^{d}}2\|K_{0}\|_{\infty}|\nabla\rho_{s}^{N}(y)|\,dy$
$=2\|K_{0}\|_{\infty}\int_{\mathbb{R}^{d}}|\nabla\rho_{s}^{N}(y)|\,dy.$ (48)
Additionally, $\rho_{s}^{N}\in\mathcal{P}\left(\mathbb{R}^{d}\right)$, $\mathbb{P}$-a.s. and then by Holder’s inequality we obtain
$\int_{\mathbb{R}^{d}}|\nabla\rho_{s}^{N}|\,dx$ $=\int_{\mathbb{R}^{d}}|\nabla\rho_{s}^{N}|\frac{\sqrt{\rho_{s}^{N}}}{\sqrt{\rho_{s}^{N}}}\,dx$
$\overset{\text{Holder}}{\leq}\left(\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{s}^{N}|^{2}}{\rho_{s}^{N}}\,dx\right)^{\frac{1}{2}}\left(\int_{\mathbb{R}^{d}}\rho_{s}^{N}\,dx\right)^{\frac{1}{2}}$
$\overset{}{=}\left(\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{s}^{N}|^{2}}{\rho_{s}^{N}}\,dx\right)^{\frac{1}{2}}.$ (49)
Now from Leibniz’s rule we deduce, for $\eta>>1$
$\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{s}^{N}|^{2}}{\rho_{s}^{N}}\,dx$ $=\int_{\mathbb{R}^{d}}\frac{\left|\nabla\left(\rho_{s}\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right|^{2}}{\rho_{s}^{N}}\,dx$
$\leq 2\int_{\mathbb{R}^{d}}\frac{\left|\nabla\rho_{s}\left(\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right|^{2}}{\rho_{s}^{N}}\,dx+2\int_{\mathbb{R}^{d}}\frac{\left|\rho_{s}\nabla\left(\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right|^{2}}{\rho_{s}^{N}}\,dx$
$\overset{}{\leq}2\int_{\mathbb{R}^{d}}\rho_{s}^{N}\frac{\left|\nabla\rho_{s}\right|^{2}}{\rho_{s}^{2}}\,dx+2\int_{\mathbb{R}^{d}}\rho_{s}^{N}\left|\nabla\ln\left(\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right|^{2}\,dx$
$\overset{}{\leq}2\eta\mathcal{H}(\rho_{s}^{N}|\rho_{s})+C_{\eta}+2\int_{\mathbb{R}^{d}}\rho_{s}^{N}\left|\nabla\ln\left(\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right|^{2}\,dx,$ (50)
where in the last inequality, we are using Donsker-Varadhan inequality (82):
[…]
$+\eta\ln{\left(\int_{\mathbb{R}^{d}}2\exp{(-2\pi|x-X_{t}|^{2})}\exp{\left(\frac{C_{1}(1+|x-X_{t}|^{2})}{\eta}\right)}\,dx\right)}$
$\overset{}{\leq}\eta\mathcal{H}(\rho_{s}^{N}|\rho_{s})+C_{\eta}.$
[…]
$\left|\left\langle S_{s}^{N},V^{N}(x-\cdot)[K\ast\rho_{s}^{N}(\cdot)-K\ast\rho_{s}^{N}(x)]\right\rangle\right|$ $\leq 2\|K_{0}\|_{\infty}\left\langle S_{s}^{N},V^{N}(x-\cdot)\right\rangle$
$\times\left(2\eta\mathcal{H}(\rho^{N}_{s}|\rho_{s})+2\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)\right)^{\frac{1}{2}}$
$\leq 1/2\rho^{N}_{s}\left(2\eta\mathcal{H}(\rho^{N}_{s}|\rho_{s})+2\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)+C_{\eta}\right)^{\frac{1}{2}},$
which implies, by $\epsilon$-Young and Holder inequalities,
$I^{2}_{t}$ $\doteq\int_{0}^{t}\int_{\mathbb{R}^{d}}\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\left\langle S_{s}^{N},V^{N}(x-\cdot)[K\ast\rho_{s}^{N}(\cdot)-K\ast\rho_{s}^{N}(x)]\right\rangle\,dx\,ds$
$\leq\int_{0}^{t}\int_{\mathbb{R}^{d}}\left|\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\right|(1/2)\rho^{N}_{s}\left(2\eta\mathcal{H}(\rho^{N}_{s}|\rho_{s})+2\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)+C_{\eta}\right)^{\frac{1}{2}}\,dx\,ds$
$\leq\int_{0}^{t}\int_{\mathbb{R}^{d}}\left|\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\right|(1/2)(\rho^{N}_{s})^{1/2}(\rho^{N}_{s})^{1/2}\left(2\eta\mathcal{H}(\rho^{N}_{s}|\rho_{s})\right)^{\frac{1}{2}}\,dx\,ds$
$+\int_{0}^{t}\int_{\mathbb{R}^{d}}\left|\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\right|(1/2)\rho^{N}_{s}\left(2\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)\right)^{\frac{1}{2}}\,dx\,ds$
$+\int_{0}^{t}\int_{\mathbb{R}^{d}}\left|\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\right|(1/2)(\rho^{N}_{s})^{1/2}(\rho^{N}_{s})^{1/2}\left(C_{\eta}\right)^{\frac{1}{2}}\,dx\,ds$
$\overset{\text{Young}}{\leq}\epsilon\int_{0}^{t}\int_{\mathbb{R}^{d}}\rho^{N}_{s}\left|\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\right|^{2}\,dx\,ds+C_{\epsilon,\eta}\int_{0}^{t}\ \mathcal{H}(\rho^{N}_{s}|\rho_{s})\,ds$
$+\sqrt{2}/2\int_{0}^{t}\left(\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)\right)^{\frac{1}{2}}\int_{\mathbb{R}^{d}}\left|\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\right|(\rho^{N}_{s})^{1/2}(\rho^{N}_{s})^{1/2}\,dx\,ds$
$+\epsilon\int_{0}^{t}\int_{\mathbb{R}^{d}}\rho^{N}_{s}\left|\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\right|^{2}\,dx\,ds+TC_{\eta}$
$\overset{\text{Holder}}{\leq}(2\epsilon+\sqrt{2}/2)\int_{0}^{t}\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)\,ds+C_{\epsilon,\eta}\int_{0}^{t}\ \mathcal{H}(\rho^{N}_{s}|\rho_{s})\,ds+TC_{\eta}.$ (51)
[…]
$I_{t}$ $\leq C_{\epsilon,K_{0},\eta}\int_{0}^{t}\mathcal{H}(\rho_{s}^{N}|\rho_{s})\,ds+(3\epsilon+\sqrt{2}/2)\int_{0}^{t}\mathcal{I}(\rho_{s}^{N}|\rho_{s})\,ds+TC_{\eta}.$ (52)
2.4 Application of Gronwall’s Lemma
In this subsection, we conclude our estimates, by applying Gronwall’s inequality.
Indeed, with the same computations as in [39], for all $\delta\in(0,\theta_{1})$ there exists a random variable $A_{0}$, finite almost surely, such that
$\sup_{t\in[0,T]}|M_{t}^{N}|\leq A_{0}N^{-\theta_{1}+\delta},$
with $\theta_{1}\doteq\left(\frac{1}{2}-\beta\Big(1+\frac{1}{d}\Big)\right)$.
[…]
$\mathcal{H}(\rho_{t\wedge\tau^{N}}^{N}|\rho_{t\wedge\tau^{N}})-\mathcal{H}(\rho_{0}^{N}|\rho_{0})$ $\leq C_{\epsilon,K_{0},\eta}\int_{0}^{t\wedge\tau^{N}}\mathcal{H}(\rho_{s}^{N}|\rho_{s})\,ds$
$\overset{}{+}(3\epsilon+\sqrt{2}/2)\int_{0}^{t\wedge\tau^{N}}\mathcal{I}(\rho_{s}^{N}|\rho_{s})\,ds$
$-\int_{0}^{t\wedge\tau^{N}}\mathcal{I}(\rho_{s}^{N}|\rho_{s})\,ds$
$+C_{\alpha,T}\frac{N^{\beta(1+\frac{2}{d}+2\alpha)}}{N}$
$+A_{0}N^{-\theta_{1}+\delta}$
$+TC_{\eta}$ (53)
and then for $\eta>>1$ and $\epsilon<<1$, we deduce
$\mathcal{H}(\rho_{t\wedge\tau^{N}}^{N}|\rho_{t\wedge\tau^{N}})-\mathcal{H}(\rho_{0}^{N}|\rho_{0})$ $\lesssim\int_{0}^{t\wedge\tau^{N}}\mathcal{H}(\rho_{s}^{N}|\rho_{s})\,ds+N^{-\theta_{2}}+A_{0}N^{-\theta_{1}+\delta}+C,$ (54)
with $\theta_{1}\doteq\left(\frac{1}{2}-\beta\Big(1+\frac{1}{d}\Big)\right)$ and $\theta_{2}\doteq 1-\beta(1+\frac{2}{d}+2\alpha)$.
Finally, from Gronwall’s Lemma in (54), recalling (28), (31) and (32), for all $\omega\in\Lambda_{0}$, there exists $N_{0}(\omega)$ such that for $N\geq N_{0}(\omega)$, $\tau^{N}(\omega)>0$ and
$\sup_{t\in[0,T]}\mathcal{H}(\rho_{t\wedge\tau^{N}}^{N}|\rho_{t\wedge\tau^{N}})$ $\lesssim\left(\mathcal{H}(\rho_{0}^{N}|\rho_{0})+N^{-\theta}+A_{0}N^{-\theta}+C\right)$ (55)
with
$\theta\doteq\min\left(1-\beta(1+\frac{2}{d}+2\alpha);\left(\frac{1}{2}-\beta\Big(1+\frac{1}{d}\Big)\right)-\delta\right).$
2.5 End of the proof: Removing the stopping time
We now remove the stopping time $\tau^{N}$ introduced in (55) by exploiting the probabilistic framework provided by Assumptions $\left(\mathbf{A}^{\rho_{0}}\right)$ and $\left(\mathbf{A}^{X_{0}}\right)$. These assumptions ensure sufficient control on the initial data and on the particle system, allowing us to pass from the localized estimates to global-in-time bounds with high probability, and finally eliminate the stopping time.
First, since $K=\nabla\cdot K_{0}$, $K_{0}\in L^{\infty}$ and $V^{N}=N^{\beta}V(N^{\beta/d}\cdot)$, by (2), we have
$\sup_{t\in[0,T]}|X_{t}^{i,N}|$ $\leq|X_{0}^{i,N}|+T\|\nabla V\|_{1}\|K_{0}\|_{\infty}N^{\beta}$
$+\sup_{t\in[0,T]}\left|\int_{0}^{t}\sqrt{2}dW_{s}^{i,N}\right|+\sup_{t\in[0,T]}\left|\int_{0}^{t}\sigma_{s}dB_{s}\right|,$ (56)
which implies for $\widetilde{C}_{K_{0},V}\doteq(1-T\|\nabla V\|_{1}\|K_{0}\|_{\infty})$
$\{\tau^{N}<T\}$ $\subset\bigcup_{i=1}^{N}\left\{\sup_{t\in[0,T]}|X_{t}^{i,N}|\geq N^{\beta}\right\}$
$=\left\{\exists i;\sup_{t\in[0,T]}|X_{t}^{i,N}|\geq N^{\beta}\right\}$
$\subset\left\{\exists i;|X_{0}^{i,N}|+\sup_{t\in[0,T]}\left|\int_{0}^{t}\sqrt{2}dW_{s}^{i,N}\right|+\sup_{t\in[0,T]}\left|\int_{0}^{t}\sigma_{s}dB_{s}\right|\geq\widetilde{C}_{K_{0},V}N^{\beta}\right\}$
$\subset L_{1}^{N}\cup L_{2}^{N}\cup L_{3}^{N},$ (57)
with
$L_{1}^{N}\doteq\left\{\exists i;|X_{0}^{i,N}|\geq C_{K_{0},V}N^{\beta}\right\},$ (58)
$L_{2}^{N}\doteq\left\{\exists i;\sup_{t\in[0,T]}\left|\int_{0}^{t}dW_{s}^{i,N}\right|\geq C_{K_{0},V}N^{\beta}\right\}$ (59)
and
$L_{3}^{N}\doteq\left\{\exists i;\sup_{t\in[0,T]}\left|\int_{0}^{t}\sigma_{s}dB_{s}\right|\geq C_{K_{0},V}N^{\beta}\right\}$ (60)
with $C_{K_{0},V}\doteq(1-T\|\nabla V\|_{1}\|K_{0}\|_{\infty})/3\sqrt{2}>0$, for $T<(1/\|\nabla V\|_{1}\|K_{0}\|_{\infty})$.
We note that, with the same computations like those led (29), we deduce
$\mathbb{P}\left(L_{1}^{N}\right)$ $\overset{}{\leq}N(C_{K_{0},V}N^{\beta})^{d-2}e^{-(C_{K_{0},V}N^{\beta})^{2}}$ (61)
We now take care about the $L_{2}^{N}$. Indeed, we note that $W_{t}^{i,N}=\int_{0}^{t}dW_{s}^{i,N}$, and since
$|W_{t}^{i,N}|^{2}=\sum_{j=1}^{d}(W_{t}^{i,N,j})^{2},$
if $|W_{t}^{i,N}|\geq a$, then there exists j=1,,dj=1,...,d, such that $|W_{t}^{i,N,j}|\geq\frac{a}{\sqrt{d}}.$ Therefore,
$\left\{\sup_{t\leq T}|W_{t}^{i,N}|\geq a\right\}\subset\bigcup_{j=1}^{d}\left\{\sup_{t\leq T}|W_{t}^{i,N,j}|\geq\frac{a}{\sqrt{d}}\right\},$ (62)
we obtain
[…]
$\leq\sqrt{\frac{T}{2\pi}}\frac{4dN}{\frac{C_{K_{0},V}N^{\beta}}{\sqrt{d}}}$
$\times\exp{\left(-\left(\frac{C_{K_{0},V}N^{\beta}}{\sqrt{d}}\right)^{2}/(2T)\right)},$ (63)
where in the second inequality, we are using (8.3)’ in Chapter 2 of [25].
Now we will derive an analogous estimate for $L_{3}^{N}$. First, we observe that, if $\sigma=0$ we are done, thus suppose $\sigma\neq 0$. Since $\sigma$ is bounded, for $j\in\{1,...,d\}$, the function defined by $t\mapsto\langle X^{j}\rangle_{t}$ is continuous and increasing. Hence, recalling (18), by Theorem 4.6, in Chapter 3 of [25], for all $j\in\{1,...,d\}$, we get
$\sup_{t\leq T}|X_{t}^{j}|=\sup_{t\leq T}\left|\widetilde{B}^{j}_{\langle X^{j}\rangle_{t}}\right|\leq\sup_{t\leq\|\sigma\|_{\infty}^{2}T}|\widetilde{B}^{j}_{t}|,$ (64)
where $(\widetilde{B}^{j}_{t})_{t\geq 0}$ is the time-change process, which in particular, is a one dimensional standard Brownian motion. In addition, since
$|X_{t}|^{2}=\sum_{j=1}^{d}|X_{t}^{j}|^{2}$
if $|X_{t}|\geq a$, then there exists j=1,,dj=1,...,d, such that $|X_{t}^{j}|\geq\frac{a}{\sqrt{d}}.$ Therefore, for all $a>0$, we find
[…]
It follows that, by (64) and (65)
[…]
$\leq\sqrt{\frac{\|\sigma\|_{\infty}^{2}T}{2\pi}}\frac{4dN}{\frac{C_{K_{0},V}N^{\beta}}{\sqrt{d}}}$
$\times\exp{\left(-\left(\frac{C_{K_{0},V}N^{\beta}}{\sqrt{d}}\right)^{2}/(2\|\sigma\|_{\infty}^{2}T)\right)},$ (66)
where in the second inequality, we are using (8.3)’ in Chapter 2 of [25].
Hence, by (61), (63) and (66) in (57), we find
$\sum_{N\in\mathbb{N}}\mathbb{P}\left(\tau^{N}<T\right)<\infty$ (67)
and then by Borel-Cantelli Lemma, we arrive at
$\mathbb{P}\left(\bigcap_{M=1}^{\infty}\bigcup_{N=M}^{\infty}\{\tau^{N}<T\}\right)=0.$ (68)
Finally, by (32) and (68), for $T<\min{\left(T_{1};(1/\|\nabla V\|_{1}\|K_{0}\|_{\infty})\right)}$ for all $\omega\in\Lambda$, with
$\Lambda\doteq\left(\bigcup_{M=1}^{\infty}\bigcap_{N=M}^{\infty}(\Lambda^{N})^{c}\right)\cap\left(\bigcup_{M=1}^{\infty}\bigcap_{N=M}^{\infty}\{\tau^{N}=T\}\right),$ (69)
there exists $N_{0}(\omega)$ such that for $N\geq N_{0}(\omega)$, $\tau^{N}(\omega)=T$ and (55) holds. It follows that, almost surely
$\sup_{t\in[0,T]}\mathcal{H}(\rho_{t}^{N}|\rho_{t})$ $\lesssim\left(\mathcal{H}(\rho_{0}^{N}|\rho_{0})+N^{-\theta}+A_{0}N^{-\theta}+C\right)$ (70)
which conclude the proof of Theorem.
3 Appendix
This appendix contains the proof of Theorem 1 together with several auxiliary inequalities used in the previous sections. We begin by establishing Theorem 1, and subsequently derive the additional estimates required throughout the paper.
3.1 Proof of Theorem 1
3.1.1 Solving (1) for $\sigma=0$
When $\sigma=0$, a solution for (1) in the sense of Theorem 1, with exception of (17), was constructed in Sections 3 and 4 of [48] and [49]. With this in view, we only need to verify that, it is possible to take $\tilde{C}_{3}$ sufficiently small, in order to get that Gaussian decay.
Proof of (17).
By Theorem 3 in [13] along with (12), we obtain there exists $C>0$ such that for all $(t,x)\in[0,T_{1}]\times\mathbb{R}^{d}$
$\bar{\rho}(t,x)\leq C\widetilde{C}_{3}\int_{\mathbb{R}^{d}}\frac{1}{t^{d/2}}\exp{\left(-\frac{\widetilde{C}_{3}|x-y|^{2}+8t|y|^{2}}{8t\widetilde{C}_{3}}\right)}dy.$ (71)
Now since
$|x-y|^{2}=|x|^{2}-2x\cdot y+|y|^{2}.$
we find
$\widetilde{C}_{3}|x-y|^{2}+8t|y|^{2}$ $=\widetilde{C}_{3}|x|^{2}-2\widetilde{C}_{3}x\cdot y+\widetilde{C}_{3}|y|^{2}+8t|y|^{2}$
$=\widetilde{C}_{3}|x|^{2}-2\widetilde{C}_{3}x\cdot y+(\widetilde{C}_{3}+8t)|y|^{2}.$ (72)
Additionally, we get
$(\widetilde{C}_{3}+8t)|y|^{2}-2\widetilde{C}_{3}x\cdot y.=(\widetilde{C}_{3}+8t)\left(|y|^{2}-2\frac{\widetilde{C}_{3}}{\widetilde{C}_{3}+8t}x\cdot y\right).$
and since
$|y|^{2}-2ax\cdot y=|y-ax|^{2}-a^{2}|x|^{2},$
with
$a=\frac{\widetilde{C}_{3}}{\widetilde{C}_{3}+8t}.$
we arrive at
$\Big((\widetilde{C}_{3}+8t)|y|^{2}-2\widetilde{C}_{3}x\cdot y\Big)=(\widetilde{C}_{3}+8t)|y-ax|^{2}-\frac{\widetilde{C}_{3}^{2}}{\widetilde{C}_{3}+8t}|x|^{2}.$ (73)
Therefore by (72) and (73) we derive
$\exp\left(-\frac{\widetilde{C}_{3}|x-y|^{2}+8t|y|^{2}}{8t\widetilde{C}_{3}}\right)=\exp{\left(-\frac{|x|^{2}}{\widetilde{C}_{3}+8t}\right)}\exp\left(-\frac{\widetilde{C}_{3}+8t}{8t\widetilde{C}_{3}}|y-ax|^{2}\right).$ (74)
and letting z=yaxz=y-ax we have
$\bar{\rho}(x,t)$ $\leq C\widetilde{C}_{3}\frac{1}{t^{d/2}}e^{-\frac{|x|^{2}}{\widetilde{C}_{3}+8t}}\int_{\mathbb{R}^{d}}\exp{\left(-\frac{\widetilde{C}_{3}+8t}{8t\widetilde{C}_{3}}|z|^{2}\right)}dz$
$\leq C\widetilde{C}_{3}\frac{1}{t^{d/2}}e^{-\frac{|x|^{2}}{\widetilde{C}_{3}+8t}}\left(\frac{\pi 8t\widetilde{C}_{3}}{\widetilde{C}_{3}+8t}\right)^{d/2}$
$\leq C\widetilde{C}_{3}(\pi 8\widetilde{C}_{3})^{d/2}\frac{1}{(\widetilde{C}_{3}+8t)^{d/2}}\exp{\left(-\frac{|x|^{2}}{\widetilde{C}_{3}+8t}\right)}$ (75)
since
$\int_{\mathbb{R}^{d}}e^{-a|z|^{2}}dz=\left(\frac{\pi}{a}\right)^{d/2}.$
It follows that taking
$\widetilde{C}_{3}<C_{3}\doteq\min\left(\frac{1}{2\pi};\frac{1}{(8\pi)^{d/2}C}\right),$ (76)
for $t\in(0,T]$ such that
$t<T_{1}\doteq\min\left(\frac{\widetilde{C}_{3}}{8};\frac{8(1-2\pi\widetilde{C}_{3})}{2\pi}\right)$ (77)
we end up with
$\bar{\rho}(x,t)\leq 2\exp{(-2\pi|x|^{2})}.$ (78)
3.1.2 Solving (1) for $\sigma\neq 0$
We assume $\tilde{\rho}$ solves the deterministic PDE
$\partial_{t}\tilde{\rho}-\Delta\tilde{\rho}+\nabla\cdot(\tilde{\rho}_{t}K\ast\tilde{\rho}_{t})=0$ (79)
Now we define $\rho(t,x)\doteq\tilde{\rho}(t,x-X_{t})$, where
$dX_{t}=\sigma_{t}dB_{t},$ (80)
Hence, by applying the Itô’s formula, since that $\sigma$ does not depends on spatial variable, we find that $\rho$ solves (1), in the sense of Theorem 1.
3.2 Inequalities
In this section, we present several inequalities used throughout the text. The first one is the classical Csiszár-Kullback-Pinsker inequality, which is a fundamental result in information theory. It quantifies how control of the relative entropy implies control of the total variation distance.
Lemma 3 (Csiszár-Kullback-Pinsker inequality).
It holds for $f,g\in\mathcal{P}\left(\mathbb{R}^{d}\right)$,
$\|f-g\|_{1}^{2}\lesssim\mathcal{H}(f|g).$ (81)
Now, we recall the Donsker–Varadhan inequality (see [23]) which follows from the variational representation of the relative entropy. It provides a tool to estimate the expectation of a test function under a non-factorized probability density and, in particular, allows for a change of measure between two probability measures.
Lemma 4 (Donsker-Varadhan Inequality).
Let $f,g\in\mathcal{P}\left(\mathbb{R}^{d}\right)$, $\eta>0$ and a function $\Phi$ on $\mathbb{R}^{d}$ such that $g\exp{\left(\Phi/\eta\right)}\in L^{1}$. Then, we have
$\int_{\mathbb{R}^{d}}f\Phi\,dx\leq\eta\mathcal{H}\left(f|g\right)+\eta\ln{\left(\int_{\mathbb{R}^{d}}g\exp{\left(\Phi/\eta\right)}\,dx\right)}.$ (82)
Acknowledgements
A. B. de Souza is partially supported by Coordenação de Aperfeiçoamento de Pessoal de Nível Superior – Brasil (CAPES) – Finance Code $001$.